# Execution-Feedback Driven Test Generation from SWE Issues

### v1

@@ 4.1. Performance of e-Otter and e-Otter++ (RQ1)
Table 1 compares e-Otter and e-Otter++ with prior approaches on both TDD-Bench Verified and SWT-bench Lite. For TDD-Bench Verified and SWT-bench Lite, the best baselines report 38.8% and 37.7%, respectively. Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively. We perform McNemar’s test (McNemar, 1947) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$). McNemar’s test compares the performance of two systems on the same set of binary-labeled instances, highlighting statistically significant differences in their error patterns. We also report our performance with GPT-4 and GPT-4o to demonstrate model agnosticity. We find that the GPT-4o-based e-Otter++ significantly outperforms all prior GPT-4 and GPT-4o-based solutions. Table 1 omits some relatively weaker approaches from both benchmarks to reduce clutter. Apart from the works listed in Table 1, the SWT-bench Leaderboard has another entry, AEGIS, but hides it by default unless a “script mode” toggle is selected. AEGIS tests are not integrated with the project’s testing framework, unlike other leaderboard entries and e-Otter++. Under that setting, AEGIS reports a 47.8% $F\!\!\to\!\!P$ rate with an undisclosed model. Integrating tests with a testing framework enables future regression testing and coverage measurements. Although not directly comparable to AEGIS, e-Otter++ with Claude-3.7-Sonnet outperforms AEGIS on SWT-bench Lite with a 52.5% $F\!\!\to\!\!P$ rate. Note that, as mentioned in Section 2, the two benchmarks have different implementations. We primarily used the TDD-Bench harness because of its speed. We also ran the SWT-Bench harness and observed a slight drop in performance (50.7%) due to dependency issues and regression tests.
While the SWT-bench paper (Mündler et al., 2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset. SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (Chowdhury et al., 2024). Although both benchmarks have some non-overlapping samples, systems usually perform similarly on them. For example, Otter++ (GPT-4o) reports 37.0% on both TDD-Bench Verified and SWT-bench Verified. Given this similarity between the two “verified” benchmarks, we opted to use only one of them to discuss our results in this paper.
@@ 4.2. Effectiveness of Test Repair (RQ2)
We just saw that that critics play an important role in test repair. But how good are LLMs as critics? We use a GPT-4o based critic on different tests and see how it performs. We take GPT-4o based Otter, Otter++, and golden (i.e., developer-written) tests from TDD-Bench Verified. We feed the test and test log and ask the model to comment on whether the test is failing for the right reason mentioned in the issue description. Table 4 shows that GPT-4o achieves a 0.78 to 0.83 F-score as a critic, which indicates the benefit of using LLMs as critics.
@@ 4.3. Impact of Heterogeneous Prompts (RQ3)
All $F\!\!\to\!\!P$ numbers we reported so far were “@ 1”, measuring the rate at which a single test is fail-to-pass. Here, we instead explore $F\!\!\to\!\!P$ @ N, i.e., the rate at which at least one among N generated tests is fail-to-pass. That is an upper bound for how well a system can do by picking a single test to submit from multiple generated candidate tests. Morphs and masks yield heterogeneous prompts, which in turn increase the diversity of candidate tests, thus hopefully increasing $F\!\!\to\!\!P$ @ N. We reuse 5 heterogeneous prompts including masks from Otter++ (Ahmed et al., 2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts. Table 5 shows that morphs increase $F\!\!\to\!\!P$ @ N, even without masks, both before (Otter) and after (e-Otter) test repair. On e-Otter, the Claude-3.7-Sonnet model with “plan + mask + morph” increases $F\!\!\to\!\!P$ @ N by 17.1% and 23.8% compared to “plan + mask” on SWT-bench Lite and TDD-Bench Verified, respectively. We observe similar improvements with the GPT-4o model on both benchmarks.
@@
Prior works (Xia et al., 2025; Ehrlich et al., 2025; Chen et al., 2021) mostly depend on high temperature for multi-sampling. Figure 5 compares the results of heterogeneous prompting with the samples generated with high temperature. We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (Chen et al., 2021). However, our approach, “plan + mask + morph”, significantly outperforms samples generated with high temperature. In Figure 5, for each $N$ (1-10), we take all possible combinations from 10 candidates and show the confidence interval in the plot.
@@ 4.5. Validating SWE-patches (RQ5)
The tests generated by our approach can be used for discarding bad code patches and increasing the precision of solutions proposed by different systems from the SWE-bench Verified leaderboard. Ahmed et al. (Ahmed et al., 2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++. They filtered out a code patch if all the tests failed on it. This achieved a precision of 65% to 92% while maintaining a decent recall of 30%-41%. Mündler et al. (Mündler et al., 2024) performed a similar experiment with one system. Note that neither of these two prior approaches could achieve a $F\!\!\to\!\!P$ rate beyond 50% even with 5 samples. Therefore, the recall suffers significantly. In contrast, e-Otter++ achieves a 71.7% $F\!\!\to\!\!P$ @ 10 rate, which could significantly benefit the recall. To showcase the efficacy of our tests in conjunction with top SWE agents, we pick the top SWE agents from the leaderboard where the generated code patches were available. Unfortunately, due to a recent change in the SWE-Bench submission process, we could not retrieve them for some systems. Table 8 shows the old precision, recall, and new precision on SWE-Bench-Verified using both Claude-3.7-Sonnet and GPT-4o models. We achieve 47.9% to 72.8% recall, regardless of the model, while substantially improving precision.
@@ 5. Discussion
5.1. Contamination through Issue Morphs
Test generation from issues could potentially suffer from model memorization and contamination since the model might have seen the data during the training phase. Both prior works by Ahmed et al. (Ahmed et al., 2025) and Mündler et al. (Mündler et al., 2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal. Since we build upon Ahmed et al.’s Otter++, we refer to their contamination analysis except for places where e-Otter++ might potentially introduce new contamination. Specifically, issue morphs might introduce contamination while rewriting the issue. To investigate this, we manually validated the samples before execution-augmented test repair. We took 50 samples (25 from each model) where the original planner prompt was not producing $F\!\!\to\!\!P$ tests but after issue morph the tests became $F\!\!\to\!\!P$. We had 4 raters with programming experience ranging from 15-30 years. We presented the raters with four pieces of information: (i) original issue description, (ii) modified issue, (iii) non $F\!\!\to\!\!P$ version of the test, (iv) $F\!\!\to\!\!P$ version of the test. We asked the raters the following three questions.
@@ 5.1. Contamination through Issue Morphs
The raters initially rated 5 samples and had a discussion on how to rate the remaining samples. After that, each rater individually rated the samples. Since we have four raters, it is infeasible to do majority voting on the rated instances. Instead, we adapted a stricter metric to assign the final label. If at least two raters think there is some addition, removal, or contamination, we counted it as “1”, otherwise “0”.
Table 9 shows the inter-rater agreement and outcome of the study. We use Fleiss’ Kappa (Fleiss, 1971) to compute inter-rater agreement, and the raters show fair to substantial agreement [^4] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement.. Note that for contamination, we observe a very low kappa value because all raters mostly choose the same category. This phenomenon is known as the kappa paradox (Feinstein and Cicchetti, 1990). Overall, 40% of modified issues have some information addition and 68% of issues have some information removal. Claude-3.7-Sonnet makes more additions and removals compared to GPT-4o. However, none of the raters think there is substantial contamination due to issue morphs. Only for 1 (2%) sample out of 50 samples, more than one rater think the rewriting might have caused some contamination. Therefore, we can infer that issue morph does not contaminate the issue. We share the ratings from each individual rater in the supplementary material.
@@
5.2. Coverage of Generated Tests
Coverage is an important metric to measure the adequacy of tests. We measure how well a generated test covers the developer-written golden code patch (note that this is not available during test generation). To do that, we find the deleted lines in $c_{\textrm{old}}$ and added lines in $c_{\textrm{new}}$ and see whether the test covers those lines. We take the total number of covered lines and divide them by the total number of deleted and added lines to compute the coverage. Figure 7 shows that $F\!\!\to\!\!P$ tests have higher coverage compared to other tests. These findings are consistent with prior works (Ahmed et al., 2025; Mündler et al., 2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
@@ 5.2. Coverage of Generated Tests
5.3. Detailed Analysis of the Tests
Table 10 presents an analysis of the e-Otter-generated tests using GPT-4o from different perspectives. We did not see much impact of focal localization on the performance, with 37.8% and 34.2% for correct and incorrect localization, respectively. Test localization has a significant impact on the performance (42.6% vs. 24.2% for correct and incorrect localization). We see that the fail-to-pass rate is higher when modify an existing test (40.2%) than when writing a new test (36.1%), which is not surprising because the model can have better context from the existing test. It is expected that the test will fail on the old codebase. However, we found 43 samples in e-Otter where the test passed on $c_{\textrm{old}}$. Also, our analysis shows that tests with assertion failures have a higher success rate (50.4% fail-to-pass rate) compared to other groups.

### v2

@@ 4.1. Performance of e-Otter and e-Otter++ (RQ1)
Table 2 compares e-Otter and e-Otter++ with prior approaches on both TDD-Bench Verified and SWT-bench Lite. For TDD-Bench Verified and SWT-bench Lite, the best baselines report 38.8% and 37.7%, respectively. Our Claude-3.7-Sonnet based e-Otter++ achieves state-of-the-art performance on both benchmarks, achieving 63.0% and 52.5% on TDD-Bench Verified and SWT-bench Lite, respectively using CodePatch + Coverage as selector (see Table 9). We perform McNemar’s test (mcnemar1947note) between our best test generator and all prior approaches and it significantly outperforms all the baselines on both TDD-Bench Verified and SWT-bench Lite with a 99% confidence interval ($p<0.01$). McNemar’s test compares the performance of two systems on the same set of binary-labeled instances, highlighting statistically significant differences in their error patterns. We also report our performance with GPT-4 and GPT-4o to demonstrate model agnosticity. We find that the GPT-4o-based e-Otter++ significantly outperforms all prior GPT-4 and GPT-4o-based solutions. Table 2 omits some relatively weaker approaches from both benchmarks to reduce clutter. Apart from the works listed in Table 2, the SWT-bench Leaderboard has another entry, AEGIS, but hides it by default unless a “script mode” toggle is selected. AEGIS tests are not integrated with the project’s testing framework, unlike other leaderboard entries and e-Otter++. Under that setting, AEGIS reports a 47.8% $F\!\!\to\!\!P$ rate with an undisclosed model. Integrating tests with a testing framework enables future regression testing and coverage measurements. Although not directly comparable to AEGIS, e-Otter++ with Claude-3.7-Sonnet outperforms AEGIS on SWT-bench Lite with a 52.5% $F\!\!\to\!\!P$ rate.
While the SWT-bench paper (mundler_et_al_2024) only describes the Lite subset of their benchmark, the leaderboard recently also added a Verified subset. SWT-bench Verified is similar to TDD-Bench Verified, as both are derived from SWE-Bench-Verified (chowdhury_et_al_2024). Although both benchmarks have some non-overlapping samples, systems usually perform similarly on them. For example, Otter++ (GPT-4o) reports 37.0% on both TDD-Bench Verified and SWT-bench Verified. Given this similarity between the two “verified” benchmarks, we opted to use only one of them to discuss our results in this paper.
@@ 4.2. Effectiveness of Test Repair (RQ2)
We just saw that that critics play an important role in test repair. But how good are LLMs as critics? We use a GPT-4o based critic on different tests and see how it performs. We take GPT-4o based Otter, Otter++, and golden (i.e., developer-written) tests from TDD-Bench Verified. We feed the test and test log and ask the model to comment on whether the test is failing for the right reason mentioned in the issue description. Table 5 shows that GPT-4o achieves a 0.78 to 0.83 F-score as a critic, which indicates the benefit of using LLMs as critics. Golden tests have no false positives, resulting in a precision of 1. However, they can have false negatives (if the model mistakenly indicates that the test is failing for the wrong reason).
@@ 4.3. Impact of Heterogeneous Prompts (RQ3)
All $F\!\!\to\!\!P$ numbers we reported so far were “@ 1”, measuring the rate at which a single test is fail-to-pass. Here, we instead explore $F\!\!\to\!\!P$ @ N, i.e., the rate at which at least one among N generated tests is fail-to-pass. That is an upper bound for how well a system can do by picking a single test to submit from multiple generated candidate tests. Morphs and masks yield heterogeneous prompts, which in turn increase the diversity of candidate tests, thus hopefully increasing $F\!\!\to\!\!P$ @ N. We reuse 5 heterogeneous prompts including masks from Otter++ (ahmed_et_al_2025) and add 5 more prompts using morphs for a total of 10 heterogeneous prompts. Table 6 shows that morphs increase $F\!\!\to\!\!P$ @ N, even without masks, both before (Otter) and after (e-Otter) test repair. On e-Otter, the Claude-3.7-Sonnet model with “plan + mask + morph” increases $F\!\!\to\!\!P$ @ N by 17.1% and 23.8% compared to “plan + mask” on SWT-bench Lite and TDD-Bench Verified, respectively. We observe similar improvements with the GPT-4o model on both benchmarks.
@@
Prior works (xia_et_al_2025; ehrlich_et_al_2025; chen2021evaluating) mostly depend on high temperature for multi-sampling. Figure 5 compares the results of heterogeneous prompting with the samples generated with high temperature. We tried three distinct temperatures (0.5, 0.8, and 1.0) and find that the best results are obtained at a temperature of 0.8, which is consistent with prior research (chen2021evaluating). However, our approach, “plan + mask + morph”, significantly outperforms samples generated with high temperature. In Figure 5, for each $N$ (1-10), we take all possible combinations from 10 candidates and show the confidence interval in the plot.
@@
Table 8 shows how much a morph can increase $F\!\!\to\!\!P$ @ 2 from $F\!\!\to\!\!P$ @ 1 (using base e-Otter) for TDD-bench-Verified because multiple morphs can independently benefit a single issue. Each morph increases $F\!\!\to\!\!P$ by 29%-32% for Claude and 11%-21% for GPT-4o.
Table 8. Contribution of each prompt to $F\!\!\to\!\!P$ @ N
Variants Claude-3.7-Sonnet GPT-4o
$F\!\!\to\!\!P$ @ 2 Change $F\!\!\to\!\!P$ @ 2 Change
Standard 235 +31.28% 194 +16.16%
Simple 235 +31.28% 186 +11.37%
dropCode 231 +29.05% 199 +19.16%
initTest 228 +27.37% 197 +17.96%
initPatch 236 +31.84% 203 +21.55%
@@ 4.5. Validating SWE-patches (RQ5)
The tests generated by our approach can be used for discarding bad code patches and increasing the precision of solutions proposed by different systems from the SWE-bench Verified leaderboard. Ahmed et al. (ahmed_et_al_2025) took the top 22 systems from the leaderboard and ran the five tests generated by Otter++. They filtered out a code patch if all the tests failed on it. This achieved a precision of 65% to 92% while maintaining a decent recall of 30%-41%. Mündler et al. (mundler_et_al_2024) performed a similar experiment with one system. Note that neither of these two prior approaches could achieve a $F\!\!\to\!\!P$ rate beyond 50% even with 5 samples. Therefore, the recall suffers significantly. In contrast, e-Otter++ achieves a 71.7% $F\!\!\to\!\!P$ @ 10 rate, which could significantly benefit the recall. To showcase the efficacy of our tests in conjunction with top SWE agents, we pick the top SWE agents from the leaderboard where the generated code patches were available. Unfortunately, due to a recent change in the SWE-Bench submission process, we could not retrieve them for some systems. Table 10 shows the old precision, recall, and new precision on SWE-Bench-Verified using both Claude-3.7-Sonnet and GPT-4o models. We achieve 47.9% to 72.8% recall, regardless of the model, while substantially improving precision.
@@
4.6. Performance on SWE-rebench (RQ6)
To show that our proposed approach is effective for non-contaminated models, we extend our evaluation on SWE-rebench with the Claude-3.7-Sonnet model. The training data cut-off date for Claude-3.7-Sonnet is November 2024 (claude-cutoff). To address data contamination, we consider instances from SWE-rebench (badertdinov2025swe) with a creation date in 2025 and check whether the developers’ written golden tests go from fail-to-pass to select the instances. This experiment uses 267 instances. Table 11 shows that execution-augmented test repair increases the fail-to-pass rate from 21% to 72% for all prompt variants. Table 12 also shows that heterogeneous prompting increases the fail-to-pass @ N by more than 32%, which is consistent with our findings on other benchmarks.
The last step of our pipeline, test selection, uses surrogate patches. We use leaderboard patches shared by the SWE-rebench authors instead of Agentless to evaluate this step. The average resolve rate of these patches is less than 40%, and these patches are also free from contamination. Table 13 shows that e-Otter++ achieved a 42.7% $F\!\!\to\!\!P$ rate on SWE-rebench, which is 75% higher than Otter. The absolute numbers on the non-contaminated SWE-rebench are notably lower than on TDD-Bench or SWT-bench, and contamination might be a major factor behind that. Another factor is that the samples from SWE-rebench come from 152 projects, most of which are less popular as the projects in SWE-bench. Popular projects are relatively older, larger, and more likely to be included in training data. Despite lower absolute numbers, our approach shows a great relative improvement, demonstrating that our two primary contributions also work for contamination-free benchmarks like SWE-rebench.
Table 11. Effectiveness of execution feedback on SWE-rebench
Prompt Otter e-Otter Change
# of F-P in % # of F-P in %
Planner 65 24.3 86 32.2 32.3%
Context Masking full 56 21.0 71 26.6 26.8%
testLoc 49 18.4 67 25.1 36.7%
patchLoc 41 15.4 60 22.5 46.3%
none 25 9.4 43 16.1 72.0%
Issue Morphing standard 60 22.5 73 27.3 21.7%
simple 45 16.9 68 25.5 51.1%
dropcode 49 18.4 75 28.1 53.1%
initTests 50 18.7 72 27.0 44.0%
initPatch 56 21.0 70 26.2 25.0%
Table 12. Impact of heterogeneous prompts on SWE-rebench
Category N Otter e-Otter
$F\!\!\to\!\!P$ @ N In % Change $F\!\!\to\!\!P$ @ N In % Change
plan + mask 5 97 36.3 NA 119 44.6 NA
plan + morph 6 115 43.1 18.6% 146 54.7 22.7%
plan + mask + morph 10 129 48.3 33.0% 158 59.2 32.8%
Table 13. Performance of e-Otter++ on SWE-rebench
Approach # of $F\!\!\to\!\!P$ in % Change from Otter
Otter 65 24.3 NA
Otter++ 87 32.6 33.8%
e-Otter 86 32.2 32.3%
e-Otter++ 114 42.7 75.4%
Finding 6. e-Otter++ is effective regardless of the data contamination issue. It improves the $F\!\!\to\!\!P$ rate by a relative 75% compared to Otter on SWE-rebench.
@@ 5. Discussion
Contamination through Issue Morphs. Test generation from issues could potentially suffer from model memorization and contamination since the model might have seen the data during the training phase. Both prior works by Ahmed et al. (ahmed_et_al_2025) and Mündler et al. (mundler_et_al_2024) addressed this issue and show that the model-generated tests are very different from developer-written tests and the impact of contamination is minimal. However, e-Otter++ might potentially introduce new contamination. Specifically, issue morphs might introduce contamination while rewriting the issue. To investigate this, we manually validated the samples before execution-augmented test repair. We took 50 samples (25 from each model) where the original planner prompt was not producing $F\!\!\to\!\!P$ tests but after issue morph the tests became $F\!\!\to\!\!P$. We had 4 raters with programming experience ranging from 15-30 years. We presented the raters with four pieces of information: (i) original issue description, (ii) modified issue, (iii) non $F\!\!\to\!\!P$ version of the test, (iv) $F\!\!\to\!\!P$ version of the test. We asked the raters the following three questions.
@@
The raters initially rated 5 samples and had a discussion on how to rate the remaining samples. After that, each rater individually rated the samples. Since we have four raters, it is infeasible to do majority voting on the rated instances. Instead, we adapted a stricter metric to assign the final label. If at least two raters think there is some addition, removal, or contamination, we counted it as “1”, otherwise “0”. Table 14 shows the inter-rater agreement and outcome of the study. We use Fleiss’ Kappa (fleiss1971measuring) to compute inter-rater agreement, and the raters show fair to substantial agreement [^3] Kappa ¡ 0: Poor, 0.01–0.20: Slight, 0.21–0.40: Fair, 0.41–0.60: Moderate, 0.61–0.80: Substantial, 0.81–1.00: Almost perfect agreement.. Note that for contamination, we observe a very low kappa value because all raters mostly choose the same category. This phenomenon is known as the kappa paradox (feinstein1990high). Overall, 40% of modified issues have some information addition and 68% of issues have some information removal. Claude-3.7-Sonnet makes more additions and removals compared to GPT-4o. However, none of the raters think there is substantial contamination due to issue morphs. Only for 1 (2%) sample out of 50 samples, more than one rater think the rewriting might have caused some contamination. Therefore, we can infer that issue morph does not contaminate the issue. We share the ratings from each individual rater in the supplementary material.
@@
Coverage of the Generated Tests. Coverage is an important metric to measure the adequacy of tests. We measure how well a generated test covers the developer-written golden code patch (note that this is not available during test generation). To do that, we find the deleted lines in $c_{\textrm{old}}$ and added lines in $c_{\textrm{new}}$ and see whether the test covers those lines. We take the total number of covered lines and divide them by the total number of deleted and added lines to compute the coverage. Figure 7 shows that $F\!\!\to\!\!P$ tests have higher coverage compared to other tests. These findings are consistent with prior works (ahmed_et_al_2025; mundler_et_al_2024) and motivate using coverage as a tie-breaker for test selection (of course, test selection uses coverage on generated code patches, not golden ones).
@@
Detailed Analysis of the Tests. Table15 presents an analysis of the e-Otter-generated tests using GPT-4o from different perspectives. We did not see much impact of focal localization on the performance, with 37.8% and 34.2% for correct and incorrect localization, respectively. Test localization has a significant impact on the performance (42.6% vs. 24.2% for correct and incorrect localization). We see that the fail-to-pass rate is higher when modify an existing test (40.2%) than when writing a new test (36.1%), which is not surprising because the model can have better context from the existing test. It is expected that the test will fail on the old codebase. However, we found 43 samples in e-Otter where the test passed on $c_{\textrm{old}}$. Also, our analysis shows that tests with assertion failures have a higher success rate (50.4% $F\!\!\to\!\!P$ rate) compared to other groups.
@@
Cost and Time. Our system, e-Otter++, uses both sequential inference scaling (via test repair) and parallel inference scaling (via heterogeneous prompting). Furthermore, we also use Agentless code patches for test selection, which requires several LLM calls. We did not generate the code patches for Agentless ourselves; instead, we relied on code patches shared by the authors for both models. However, based on their reported numbers and our experiments, we can estimate the cost for each sample. For the Claude and GPT-4o models, our estimated cost is \$2.75 and \$1.80 per instance, respectively. Note that prior test generation approaches do not use code patches for test selection, which adds more cost. However, it is likely that a reproduction test generation approach will be paired with a SWE-patch generation approach. In that case, the additional cost of generating code patches will be amortized. Additionally, excluding Agentless, e-Otter++ costs \$1.70 (\$0.17/candidate) for Claude and \$1.10 (\$0.11/candidate) for GPT-4o. Figure 5 shows how the $F\!\!\to\!\!P$ @ N changes with the number of candidates. For example, 6 samples yield 230 $F\!\!\to\!\!P$ tests and generating 6 samples costs \$0.66. It is hard to report the end-to-end time because several steps of our process can be parallelized. We did most of our experiments sequentially due to shared resources. Docker setup takes 5-10 minutes based on the repository. Other steps (i.e., test variant generation, agentless generation, ranker) can take up to 2–3 minutes, some of which are parallelizable.